OpenStudy (rane):

Alvaro, an archer, aims an arrow at a 25 angle at an apple across field. Alvaro fires the arrows from a height of 1.0m with an initial speed of 33m/s and hits the apple, which is 4m above the ground. how far away ( horizontal distance, X) is Alvaro fro the apple?

OpenStudy (unklerhaukus):



OpenStudy (rane):

i have drew the same diagram, i just wana know how to solve it

OpenStudy (unklerhaukus):

is there some information that is missing?

OpenStudy (rane):

thats all wht I'm given

OpenStudy (unklerhaukus):

maybe it is more like this 

OpenStudy (unklerhaukus):

\[y(t)=y_0+v_{y_0}\cdot t+\tfrac12a_y\cdot t^2\\\quad=y_0+v_0\cdot\sin\theta\cdot t-\tfrac12g\cdot t^2\\ \qquad=1+33\cdot\sin(25°)\cdot t-\tfrac12\cdot9.8\cdot t^2\] \[y(t_f)=1+33\cdot\sin(25°)\cdot t_f-\tfrac12\cdot9.8\cdot t_f^2=4\]

OpenStudy (unklerhaukus):

\[x(t)=x_0+v_{x_0}\cdot t+\tfrac12a_x\cdot t^2\\ \quad =v_0\cdot\cos\theta\cdot t\\ \qquad=33\cdot\cos(25°)\cdot t \] \[x(t_f)=33\cdot\cos(25°)\cdot t_f\]

OpenStudy (rane):

can u pls explain it first, i don't want answers i wana know how to solve it my own